# Rajamanickam

### v4

@@ Abstract
Strong anchoring boundary conditions are conventionally modelled by imposing Dirichlet conditions on the order parameter in Landau–de Gennes theory, neglecting the finite surface energy of realistic anchoring. This work revisits the strong anchoring limit for nematic liquid crystals in confined two-dimensional domains. By explicitly retaining a Rapini–Papoular surface energy and adopting a scaling where the extrapolation length $l_{ex}$ is comparable to the coherence length $\xi$, we analyse both the small-domain ($\varepsilon=h/\xi\to 0$; $h$ is the domain size) and Oseen–Frank $(\varepsilon\to\infty$) asymptotic regimes. In the small-domain limit, the leading-order equilibrium solution is given by the average of the boundary data, which can vanish in symmetrically frustrated geometries, leading to isotropic melting. In the large-domain limit, matched asymptotic expansions reveal that surface anchoring introduces an $O(1/\varepsilon)$ correction to the director field near boundaries, in contrast to the $O(1/\varepsilon^{2})$ correction predicted by Dirichlet conditions. The analysis captures the detailed structure of interior and boundary defects, showing that mixed (Robin-type) boundary conditions yield smoother defect cores and more physical predictions than rigid Dirichlet conditions. Numerical solutions for square and circular wells with tangential anchoring illustrate the differences between the two boundary condition treatments, particularly in defect morphology. The results demonstrate that a consistent treatment of anchoring energetics is essential for accurate modelling of nematic equilibria in micro- and nano-scale confined geometries.
@@ 3 Problem formulation in the reduced two-dimensional Landau–de Gennes framework
with $\boldsymbol{\nu}$ being the unit outward normal to $\partial\Omega$.
@@ 5.1 Outer solution
$\mathbf{n}=\mathbf{n}_{0}+\varepsilon\varphi_{1}\mathbf{t}_{0}+\varepsilon^{2}(\varphi_{2}\mathbf{t}_{0}-\tfrac{1}{2}\varphi_{1}^{2}\mathbf{n}_{0})+\cdots$ (42)
@@
with $\mathbf{Q}_{0}:\mathbf{T}_{0}=0$. In the limit $\gamma\to\infty$ (i.e., Dirichlet boundary conditions), $\varphi_{1}=0$. For this case, $O(\varepsilon^{-2})$ term proportional to $\mathbf{Q}_{0}$ was obtained earlier by Di Fratta et. al. [di2020landau, Section 4] as an $O(\varepsilon^{-2})$ correction affecting only the scalar order parameter. Our expansion provides the full $O(\varepsilon^{-2})$ correction in $\mathbf{Q}$, including the component $\varphi_{2}\mathbf{T}_{0}$ orthogonal to $\mathbf{Q}_{0}$, which corresponds to a second-order correction in the director field not previously computed.
@@ 6 Numerical results with tangential anchoring
In this section, we present numerical solutions of the problem (12)-(14) for few illustrative domains subjected to tangential anchoring, with $s_{b}=2(q_{1,b}^{2}+q_{2,b}^{2})^{\frac{1}{2}}=1$. The stability of the computed solutions is also evaluated; the details of the stability analysis are provided in Appendix A. In all numerical computations, we set $\gamma=1$. For comparison, we also consider the $\varepsilon$-independent Dirichlet boundary condition
@@ Acknowledgment
with $|\hat{\mathbf{q}}|\ll|\bar{\mathbf{q}}|$ and linearising the gradient-flow equations, we obtain the eigenvalue problem [rajamanickam2026nematic]
@@
The bulk potential terms in the Hessian are diagonal in this basis. Since $|\bar{\mathbf{q}}|^{2}\in[0,1/4]$, we have $12|\bar{\mathbf{q}}|^{2}-1\in[-1,2]$ and $4|\bar{\mathbf{q}}|^{2}-1\in[-1,0]$. The parallel term $12|\bar{\mathbf{q}}|^{2}-1$ thus stabilises magnitude fluctuations when $\bar{s}$ is close to unity, but can destabilize in low-order regions such as defect cores. The perpendicular term $12|\bar{\mathbf{q}}|^{2}-1$ governs how director rotations couple to the local degree of order. In perfectly ordered regions, where $\bar{s}=1$, this term is zero, it vanishes, reflecting rotational symmetry and producing soft elastic modes. In regions of reduced order, it is negative, meaning the bulk potential provides no restoring force for director rotations, which then rely primarily on elasticity for stability. Together, these effects imply that defect cores are inherently vulnerable to instabilities in both magnitude and orientation, a phenomenon absent in the director‑only Oseen–Frank approximation.
@@
$\nabla^{2}\hat{\varphi}+(\tfrac{1}{2}\sigma-4|\nabla\bar{\varphi}|^{2})\hat{\varphi}=0,\qquad\hat{\varphi}_{\partial\Omega}=0.$ (90)
This eigenvalue problem is formally analogous to the two-dimensional Schrödinger equation
$\nabla^{2}\psi+\frac{2m}{\hbar^{2}}(E-V)\psi=0$ (91)
@@
Since the ”potential” $4|\nabla\bar{\varphi}|^{2}\geq 0$, it follows that the ”energy” $\sigma/2\geq 0$, indicating that the base state is linearly stable. This reflects the strength and limitation of the Oseen–Frank approximation because elastic stability of the director field is captured while instabilities associated with variations in the magnitude of $\mathbf{Q}$ are neglected. In the full Landau–de Gennes theory, variations in $s$ can produce negative eigenvalues corresponding to core instabilities of defects. Within such cores, both $\hat{q}_{\parallel}$ and $\hat{q}_{\perp}$ must be retained, and the above analysis provides only an approximate description.

### v5

@@ Abstract
Strong anchoring boundary conditions are conventionally modelled by imposing Dirichlet conditions on the order parameter in Landau–de Gennes theory, neglecting the finite surface energy of realistic anchoring. This work revisits the strong anchoring limit for nematic liquid crystals in confined two-dimensional domains. By explicitly retaining a Rapini–Papoular surface energy and adopting a scaling where the extrapolation length $l_{ex}$ is comparable to the coherence length $\xi$, we analyse both the small-domain ($\varepsilon=h/\xi\to 0$; $h$ is the domain size) and Oseen–Frank $(\varepsilon\to\infty$) asymptotic regimes. In the small-domain limit, the leading-order equilibrium solution is given by the average of the boundary data, which can vanish in symmetrically frustrated geometries, leading to isotropic melting. In the large-domain limit, matched asymptotic expansions reveal that surface anchoring introduces an $O(1/\varepsilon)$ correction to the director field near boundaries, in contrast to the $O(1/\varepsilon^{2})$ correction predicted by Dirichlet conditions. The analysis captures the detailed structure of interior and boundary defects, showing that mixed (Robin-type) boundary conditions yield smoother defect cores and more physical predictions than rigid Dirichlet conditions. Importantly, a stability analysis based on the intrinsic-basis decomposition shows that even small deviations of $O(1/\varepsilon^{2})$ from perfect scalar order can destabilise some Oseen–Frank solutions, indicating that only a restricted subset of solutions are realised as stable equilibria in the full Landau–de Gennes framework. Numerical solutions for square and circular wells with tangential anchoring illustrate the differences between the two boundary condition treatments, particularly in defect morphology. These results demonstrate that a consistent treatment of anchoring energetics, together with stability considerations, is essential for accurate modelling of nematic equilibria in micro- and nano-scale confined geometries.
@@ 3 Problem formulation in the reduced two-dimensional Landau–de Gennes framework
with $\boldsymbol{\nu}$ being the unit outward normal to $\partial\Omega$. The formulation for the stability analysis is provided in Appendi x A.
@@ 5.1 Outer solution
$\mathbf{n}=\mathbf{n}_{0}+\varepsilon^{-1}\varphi_{1}\mathbf{t}_{0}+\varepsilon^{-2}(\varphi_{2}\mathbf{t}_{0}-\tfrac{1}{2}\varphi_{1}^{2}\mathbf{n}_{0})+\cdots$ (42)
@@
with $\mathbf{Q}_{0}:\mathbf{T}_{0}=0$. The $\varepsilon^{-2}\mathbf{Q}_{0}$ in (44) corresponds to ${\partial^{2}\mathbf{Q}}/{\partial\varphi^{2}}=-4\mathbf{Q}$, whereas the $\varepsilon^{-2}\mathbf{T}_{0}$ corresponds to the mixed derivative ${\partial^{2}\mathbf{Q}}/{\partial s\partial\varphi}=\mathbf{T}$, evaluated at $s=1$ and $\varphi=\varphi_{0}$. In fact, the exact closed-form expansion for the $\mathbf{Q}(s,\varphi)$ about $s=1$ and $\varphi=\varphi_{0}$ is
$\mathbf{Q}(s,\varphi)=(1+\delta s)\left[\cos(2\delta\varphi)\mathbf{Q}_{0}+\tfrac{1}{2}\sin(2\delta\varphi)\mathbf{T}_{0}\right],\quad\delta s=s-1,\quad\delta\varphi=\varphi-\varphi_{0}$ (47)
from which the asymptotic series in $\varepsilon^{-1}$ follows upon substituting $\delta s=\varepsilon^{-2}s_{2}+\cdots$ and $\delta\varphi=\varepsilon^{-1}\varphi_{1}+\varepsilon^{-2}\varphi_{2}+\cdots$.
In the limit $\gamma\to\infty$ (i.e., Dirichlet boundary conditions), $\varphi_{1}=0$. For this case, $O(\varepsilon^{-2})$ term proportional to $\mathbf{Q}_{0}$ was obtained earlier by Di Fratta et. al. [di2020landau, Section 4] as an $O(\varepsilon^{-2})$ correction affecting only the scalar order parameter. Our expansion provides the full $O(\varepsilon^{-2})$ correction in $\mathbf{Q}$, including the component $\varphi_{2}\mathbf{T}_{0}$ orthogonal to $\mathbf{Q}_{0}$, which corresponds to a second-order correction in the director field not previously computed.
@@ 5.3 Structure of a boundary defect
5.4 Remarks on the stability
Although the reduced two-dimensional Landau–de Gennes framework formally contains the Oseen–Frank theory as the limiting case $s_{0}\to 1$, this limit is structurally delicate from the perspective of stability. As shown in Appendix A, linear perturbations about a leading-order Oseen–Frank equilibrium naturally decompose into modes associated with variations in the magnitude of $\mathbf{Q}$ and rotations of the director field. In the strict Oseen–Frank limit, director perturbations are elastically stabilised, and the corresponding equilibria are always local minima with respect to director variations.
However, even a small departure from perfect order, $1-s\sim O(\varepsilon^{2})$, removes the bulk restoring force for director rotations and permits coupling to magnitude fluctuations. This is precisely captured by the higher-order corrections calculated from the Euler–Lagrange equations, where $s=1-2\varepsilon^{-2}|\nabla\varphi_{0}|^{2}+\cdots$ and $\mathbf{n}=\mathbf{n}_{0}+\varepsilon^{-1}\varphi_{1}\mathbf{t}_{0}+\varepsilon^{-2}(\varphi_{2}\mathbf{t}_{0}-\tfrac{1}{2}\varphi_{1}^{2}\mathbf{n}_{0})+\cdots$. Consequently, the higher-order terms are not merely refinements; they determine which Oseen–Frank solutions survive as stable equilibria of the full Landau–de Gennes energy and highlight the special sensitivity of defect cores and regions of reduced order. The intrinsic-basis stability analysis in Appendix A makes this restriction precise and demonstrates how coupling between magnitude and director fluctuations may govern the selection of admissible solutions. This does not imply that all deviations from $s=1$ lead to instability; rather, it implies that only a restricted class of Oseen–Frank solutions can persist as stable equilibria of the full Landau–de Gennes energy.
@@ 6 Numerical results with tangential anchoring
In this section, we present numerical solutions of the problem (12)-(14) for few illustrative domains subjected to tangential anchoring, with $s_{b}=2(q_{1,b}^{2}+q_{2,b}^{2})^{\frac{1}{2}}=1$. In all numerical computations, we set $\gamma=1$. For comparison, we also consider the $\varepsilon$-independent Dirichlet boundary condition
@@ Acknowledgment
with $|\hat{\mathbf{q}}|\ll|\bar{\mathbf{q}}|$ and linearising the gradient-flow equations, we obtain the eigenvalue problem
@@
The bulk potential terms in the Hessian are diagonal in this basis. Since $|\bar{\mathbf{q}}|^{2}\in[0,1/4]$, we have $12|\bar{\mathbf{q}}|^{2}-1\in[-1,2]$ and $4|\bar{\mathbf{q}}|^{2}-1\in[-1,0]$. The parallel term $12|\bar{\mathbf{q}}|^{2}-1$ thus stabilises magnitude fluctuations when $\bar{s}$ is close to unity, but can destabilize in low-order regions such as defect cores. The perpendicular term $4|\bar{\mathbf{q}}|^{2}-1$ governs how director rotations couple to the local degree of order. In perfectly ordered regions, where $\bar{s}=1$, this term is zero, it vanishes, reflecting rotational symmetry and producing soft elastic modes. However, even a slight violation of this condition, $1-\bar{s}\sim O(\varepsilon^{2})$, renders this term negative, so that the bulk potential provides no restoring force for director rotations, which must then rely primarily on elasticity for stability. Together, these effects imply that defect cores are inherently vulnerable to instabilities in both magnitude and orientation, a phenomenon absent in the director‑only Oseen–Frank approximation.
@@
$\nabla^{2}\hat{\varphi}+(\tfrac{1}{2}\sigma-4|\nabla\bar{\varphi}|^{2})\hat{\varphi}=0,\qquad\hat{\varphi}|_{\partial\Omega}=0.$ (91)
The term $4|\nabla\bar{\varphi}|^{2}$ in the Hessian arises from the intrinsic curvature of the uniaxial constraint manifold and therefore persists in the Oseen–Frank limit, even though it does not appear in the leading-order Euler–Lagrange equations. This eigenvalue problem is formally analogous to the two-dimensional Schrödinger equation with an infinite barrier on $\partial\Omega$, i.e.,
$\nabla^{2}\psi+\frac{2m}{\hbar^{2}}(E-V)\psi=0,\qquad\psi|_{\partial\Omega}=0,$ (92)
@@
Since the “potential” $4|\nabla\bar{\varphi}|^{2}\geq 0$, it follows that the “energy” $\sigma/2\geq 0$, indicating that the base state is linearly stable. In fact, the spectrum is discrete and can be ordered as $0<\sigma_{1}<\sigma_{2}\leq\sigma_{3}\leq\cdots$ with $\sigma_{n}\geq\lambda_{n}$, where $\nabla^{2}\phi_{n}=\tfrac{1}{2}\lambda_{n}\phi_{n}$, $\phi_{n}|_{\partial\Omega}=0$. This reflects the strength and limitation of the Oseen–Frank approximation because elastic stability of the director field is captured while instabilities associated with variations in the magnitude of $\mathbf{Q}$ are neglected.
In the full Landau–de Gennes theory, however, variations in the scalar order parameter $s$ can generate negative eigenvalues associated with defect–core instabilities. Within such cores, both $\hat{q}_{\parallel}$ and $\hat{q}_{\perp}$ must be retained, and the reduced Oseen–Frank description provides only an approximate account of the stability landscape. Consequently, Oseen–Frank equilibria are always local minima with respect to director variations, but need not be minima of the full Landau–de Gennes energy.
More importantly, the intrinsic-basis analysis above shows that even a small departure from the Oseen–Frank manifold, $1-s\sim O(\varepsilon^{2})$, removes the bulk restoring force for director rotations and thereby permits destabilisation through coupling to magnitude fluctuations. This does not imply that every such deviation is unstable; rather, it implies that only a restricted class of Oseen–Frank solutions can persist as genuine equilibria of the full Landau–de Gennes theory, namely those for which this destabilising mechanism is exactly balanced by elastic and geometric effects.